part of $20000 is to be invested at 9% and the remainder is to be invested at 12%. What is the least amount of money that can be invested at 12% in order to have a yearly income of at least $2250 from the two investments?
Answer by stanbon(57984)

part of $20000 is to be invested at 9% and the remainder is to be invested at 12%. What is the least amount of money that can be invested at 12% in order to have a yearly income of at least $2250 from the two investments?
-------------------------------------
0.09x + 0.12(20,000-x) = 2250
---
9x + 12*20,000 - 12x = 225000
-3x = 225000-240,000
-3x = -15000
x = $5000
20,000-x = $15,000 (amt invested at 12%)
=====================
Cheers,
Stan H.